# Cryogenic cesium iodide as a potential PET material

### v3

@@ 3.1 Light yield and energy resolution
The detection efficiency has been estimated with a Geant4 [16] Monte Carlo simulation using the LUT optical model [17] at 58%. This relatively low efficiency is caused by the narrow form factor of the crystals (almost 7 times longer than wide), since the scintillation photons are reflected inside the crystals several times before reaching the photosensor. Thus, even if the PTFE reflectivity is approximately 95% [18], the probability for a scintillation photon of being absorbed before reaching the MPPC compounds at every reflection, becoming non-negligible.
@@
The voltage breakdown of MPPCs is proportional to the temperature with a slope of approximately 54 mV/K [13]. At K, we adjusted the bias voltage by reducing it by 10.3 V to compensate for this effect. We then repeated the single photon measurement. The integration window was increased to 5.5 , due to the longer decay time of CsI at low temperatures (see section 3.2). The MPPC dark count rate is halved every 5.3 K [19], becoming negligible for our integration time at K. Under these low-temperature conditions, the number of photoelectrons increased by over an order of magnitude. This corresponds to light yields of photons/MeV and photons/MeV for the respective crystals. The response of the photosensor was verified to be linear by triggering only one channel and measuring the amount of charge at the 1275 keV peak.
@@ 3.2 Emission time profile and coincidence time resolution
The CTR for events near the 511 keV photopeak was measured in our setup using a baseline-corrected leading edge time pick-off method [23]. The threshold was set at 5 times the electronic noise, defined as the standard deviation of the voltage values in a region with no signal. The events were selected by requiring a charge equivalent to the energy window . An example of a triggered event with an energy near the photopeak is shown in figure 4.
@@
At this temperature, the measured CTR was ns, which is qualitatively consistent with the value extrapolated from the room temperature measurement using eq. (1.1). Thus, the effect of the increase in decay time is partially compensated by an increase in the number of detected photons, mitigating the impact on the CTR. The CTR as a function of temperature and the time difference distribution at 104 K are shown in figure 6. Similarly to the case of the energy resolution, the value of CTR could be improved by adopting photosensors with a larger PDE. Following naively eq. (1.1), a CTR of 1.3 ns could be achieved at cryogenic temperatures with a PDE approximately better.
@@ 4 Summary and future prospects
The coincidence time resolution, while sufficient for PET applications, is too large to support time-of-flight (TOF) measurements. Thus, this material might be a promising candidate for TBPET scanners, where the scintillator cost represents a large fraction of the total cost and can potentially offset the added complexity of a cryogenic system.

### v4

@@ 3.1 Light yield and energy resolution
The detection efficiency has been estimated with a Geant4 [16] Monte Carlo simulation using the LUT optical model [17] at 58%. This relatively low efficiency is caused by the narrow form factor of the crystals (almost 7 times longer than wide), since the scintillation photons are reflected inside the crystals several times before reaching the photosensor. Thus, even if the PTFE reflectivity is approximately 95% [18], the probability for a scintillation photon of being absorbed before reaching the MPPC compounds at every reflection, becoming non-negligible. Different optical models have also been explored (UNIFIED [19], DAVIS [20]) obtaining variations in the value of smaller than 10%.
@@
The voltage breakdown of MPPCs is proportional to the temperature with a slope of approximately 54 mV/K [13]. At K, we adjusted the bias voltage by reducing it by 10.3 V to compensate for this effect. We then repeated the single photon measurement. The integration window was increased to 5.5 , due to the longer decay time of CsI at low temperatures (see section 3.2). The MPPC dark count rate is halved every 5.3 K [21], becoming negligible for our integration time at K. Under these low-temperature conditions, the number of photoelectrons increased by over an order of magnitude. This corresponds to light yields of photons/MeV and photons/MeV for the respective crystals.
Such a high light yield might affect the linearity of the response in a SiPM, as the probability for a photon to hit the same microcell during its recovery time is non-negligible. However, in the case of cryogenic CsI, the time emission constant is approximately 800 ns (see section 3.2), which spreads the photon emission over a relatively wide time window. Additionally, we are using SiPMs with 14,400 microcells (cell size is 25 ). A conservative assumption on the cell recovery time is 100 ns ( larger than the value obtained in ref. [22]). Thus, with a light yield of 100,000 photons/MeV and a detector efficiency of 58%, the photons that reach the sensor in the first 100 ns for an annihilation event are:
With a PDE of 17%, this corresponds to a number of fired cells of approximately 600 and an occupancy of 4.2%, which is far from the saturation regime. In addition, the response of the photosensor was verified to be linear by triggering only one channel: the charge corresponding to the Na-22 1275 keV peak is only slightly less than what one would extract by scaling linearly the charge at 511 keV, as shown in figure 3.
@@ 3.2 Emission time profile and coincidence time resolution
The CTR for events near the 511 keV photopeak was measured in our setup using a baseline-corrected leading edge time pick-off method [26]. The threshold was set at 5 times the electronic noise, defined as the standard deviation of the voltage values in a region with no signal and corresponding to approximately 3 mV. No improvement was observed with smaller threshold values and a worsening of the CTR was observed with a threshold lower than . The events were selected by requiring a charge equivalent to the energy window . An example of a triggered event with an energy near the photopeak is shown in figure 5.
@@
At this temperature, the measured CTR was ns, which is qualitatively consistent with the value extrapolated from the room temperature measurement using eq. (1.1). Thus, the effect of the increase in decay time is partially compensated by an increase in the number of detected photons, mitigating the impact on the CTR. The CTR as a function of temperature and the time difference distribution at 104 K are shown in figure 7. Similarly to the case of the energy resolution, the value of CTR could be improved by adopting photosensors with a larger PDE, already available on the market. Following naively eq. (1.1), a CTR of 1.3 ns could be achieved at cryogenic temperatures with a PDE approximately better.
Further improvements might be possible by detecting Cherenkov photons, which are emitted promptly. In this context, CsI appears promising due to its low self-absorption and relatively high refractive index in the UV range (1.9 at 350 nm), where Cherenkov emission is more abundant. A preliminary simulation – assuming a 75 cell MPPC (corresponding to a higher PDE) and a perfect time response from the electronics – indicates that a CTR of approximately 400 ps could be achieved, potentially enabling time-of-flight (TOF) measurements.
@@ 4 Summary and future prospects
The coincidence time resolution is sufficient for PET applications and might support TOF measurements with the detection of Cherenkov photons. Thus, this material might be a promising candidate for TBPET scanners, where the scintillator cost represents a large fraction of the total cost and can potentially offset the added complexity of a cryogenic system.